the rod has a diameter of 15mm and 30 cm long. the thickness of the gold layer should be 0.2mm. the electromechanical equivalent of gold is 1.118 mg and density of gold is19300 kg/m^3. if the current process is executed with a current of 130A how long will it take to complete


1
Expert's answer
2021-06-06T17:28:02-0400

"\frac{Energy }{Volume}=\rho gh = 19300*9.81*30/100=567999"

"Volume =\pi *(15/2000)^2*30/100=0.00005"

Energy = 567999*0.00005=30.1211 J

"Power =\frac{J}{s} \implies \frac{30.11211}{t}=1.118*130"

"t = \frac{1.118*130}{30.1121}= 4.8266 = 5 s"